# FlowDubber: Movie Dubbing with LLM-based Semantic-aware Learning and Flow Matching based Voice Enhancing

### v1

@@ 3.3. Dual Contrastive Aligning for Dubbing
This module is designed to solve alignment problems in dubbing by introducing a Dual Contrastive Learning (DAL) between lip movement sequence and phoneme sequence.
Lip-motion Feature Extractor. To ensure fairness for measuring the alignment ability of DAL, we first use the same extractor (Cong et al., 2023; Zhang et al., 2024, 2025a) to obtain lip motion features from silent videos $V_{s}$:
@@
where $z_{m}\in\mathbb{R}^{{L_{v}}\times{d_{m}}}$ denotes the output lip motion embedding, $L_{v}$ indicates the length of lip sequence, and $d_{m}$ is embedding size. The $\mathrm{LipCrop}(\cdot)$ uses the face landmarks tool to crop mouth area and $\mathrm{LipExtra}(\cdot)$ consists of 3D convolution, ResNet-18, and 1D convolution (Cong et al., 2023) to capture dynamic lip-motion representation.
Dual Contrastive Learning. We focus on learning the intrinsic correlation between phoneme-level pronunciation and lip movement to achieve reasonable alignment for movie dubbing. Following the contrastive learning manner, we introduce the InfoNCE loss (Tian, 2022) to encourage the model to distinguish correct lip-phoneme pairs. Specifically, we first treat the lip motion features $z_{m}$ as queries and the phoneme embeddings $z_{p}$ as keys. To establish positive pairs, we align each lip motion frame with its corresponding phoneme based on ground-truth timing annotations by MFA and FPS. This ensures that each $z_{m}^{i}$ should be maximally similar to its temporally aligned $z_{p}^{j}$, while being distinct from other phonemes:
@@
Unlike single-directional contrastive learning, which only aligns one modality to the other, the proposed DCA ensures mutual alignment, reducing ambiguities where similar phonemes might be confused. Finally, we simply use their average as dual contrastive loss:
@@
Aligning Phoneme Level Feature. The similarity matrix between textual phoneme embedding and lip movement embedding ${Sim}(z_{m},z_{p})$ is constrained by dual contrastive learning, then the ${Sim}(z_{m},z_{p})$ further guidance to the generation of aligning sequences, including: (1) lip-related aligning sequences $C_{lip}$. (2) phoneme-related aligning sequences. Specifically, the $C_{lip}$ is obtained by multi-head attention module in  (Cong et al., 2023), in which the $z_{p}$ serves as key and value, and the $z_{m}$ is the query. Unlike (Cong et al., 2023), the learnable ${Sim}(z_{m},z_{p})$ is used as multi-head attention weight matrix to provide correct relevance. Next, by monotonic alignment search (MAS) (Kim et al., 2020), the ${Sim}(z_{m},z_{p})\in\mathbb{R}^{{L_{v}}\times{L_{t}}}$ is flat to mapping table $tab\in\mathbb{R}^{{L_{t}}\times{1}}$, which records the number of video frames corresponding to each phoneme unit. Finally, the $tab$, $LLM_{p}$, $z_{p}$, and $C_{lip}$ are associated to mel-spectrograms level prior conditions $\mu$:
@@
where $\mathrm{Up}(\cdot)$ is used to expand $LLM_{p}$ and $z_{p}$ to video level according to mapping $tab$. The $\mathcal{F}(\cdot)$ indicates the fusion module, which consists of two 2D upsampling convolutional layers and transformer-based mel-decoder. The output $\mu\in\mathbb{R}^{{L_{m}}\times{d_{m}}}$, where ${l_{m}}$ and $d_{m}$ represent the length and embedding size of mel-spectrogram.
@@ 3.4. Flow-based Voice Enhancing
In this section, we introduce flow-based voice enhancing, including Style Flow Matching Prediction to inject speaker style into flow matching and LLM-based Acoustics Flow Matching Guidance to improve the clarity of generated speech by enhancing semantic information from the LLM.
Style Flow Matching Prediction. Flow matching generates mel-spectrograms $\hat{M}$ from Gaussian noise by a vector field. Given mel-spectrogram space with data $M$, where $M\sim q(M)$. We aim to train a flow matching network to fit $q(M)$ by predicting the probability density path given the vector field, which can be defined as $p_{t}(x)$. Here $t\in[0,1]$, $p_{0}(x)=\mathcal{N}(x;0,I)$ and $p_{1}(x)=q(x)$. Flow matching can predict the probability density path, gradually transforming $x_{0}\sim p_{0}(x)$ into $M\sim q(M)$. Our Flow matching prediction network is based on optimal-transport conditional flow matching (OT-CFM). OT-CFM uses a linear interpolation flow $\phi_{t}(x)=(1-(1-\sigma_{\min})t)x_{0}+tM$, which satisfies the marginal condition $\phi_{0}(x)=x_{0}$ and $\phi_{1}(x)=M$. The gradient field vector field of OT-CFM is $u_{t}(\phi_{t}(x)|M)=M-(1-\sigma_{min})x_{0}$. The training objective of Flow matching prediction network is to predict the gradient vector field $v_{t}(\phi_{t}(x)|\mu_{SATL},\theta)$, which should be close to $u_{t}(\phi_{t}(x)|M)$: Here $\mu_{SATL}$ is style-enhanced mel-spectrograms level prior according to $\mu$ in Eq. 10. To enhance speakers’ style, we introduced SATL in flow matching. Specifically, during the flow matching generation process, SATL introduces and enhances style information through affine transformation, which can be formulated as:
(11) $\mu_{SATL}=\gamma_{2}(\gamma_{1}\mu+\beta_{1})+\beta_{2},$
where $\gamma_{1},\gamma_{2},\beta_{1},\beta_{2}$ are parameters predicted by SATL based on style features. We train Style Flow Matching Prediction Network using condition $\mu_{SATL}$. We aim for the Flow Matching prediction network to generate the target mel-spectrogram M conditioned on a given $\mu_{SATL}$. During the inference process, Flow Matching prediction network solves the ODE $d\phi_{t}(x)=v_{t}(\phi_{t}(x)|\mu_{SATL},\theta)dt$ from $t=0$ to $t=1$ to generate a mel-spectrogram $\hat{M}$.
LLM-based Acoustics Flow Matching Guidance. To enhance the clarity of the generated audios, we enhanced the mel-spectrograms level prior conditions by LLM-based Acoustics Flow Matching Guidance. We observed that the generation process in LLM includes semantic tokens and text tokens, which introduce semantic knowledge. Specifically, we enhance LLM’s information in flow matching process to improve speech clarity based on classifier-free guidance, which can be formulated as:
$\tilde{v}_{t}(\phi_{t}(x)|\mu,\theta)=v_{t}(\phi_{t}(x)|\mu_{SATL},\theta)+\alpha\Big{(}v_{t}(\phi_{t}(x)|\mu_{SATL},\theta)-v_{t}(\phi_{t}(x)|\mu^{\prime},\theta)\Big{)}$
Here $\mu^{\prime}=\mathcal{F}([C_{lip},\mathrm{Up}(\phi,z_{p},tab))$, $\phi$ refers to zero vector. By enhancing only the LLM information to improve speech clarity with classifier-free guidance, we can control the mel-spectrograms clarity by removing noise and boosting overall quality. As a result, the proposed guidance mechanism strengthens the semantic information accessible to the flow-matching prediction network, thereby refining the gradient vector field generation process to achieve higher speech clarity.
Table 1. Compared with related Dubbing methods on Chem benchmark. For the Dub 1.0 setting, we use the ground truth audio as reference audio, for the Dub 2.0 setting, we use the non-ground truth audio from the same speaker within the dataset as the reference audio which is more aligned with practical usage in dubbing. M2CI-Dub (Zhao et al., 2025) has no provide setting2 audio.
@@
Imagin (Lee et al., 2023a) (ICASSP 2023) 1.98 12.50 0.250 62.24 2.62 1.96 12.53 0.184 68.13 2.13
V2C-Net (Chen et al., 2022) (CVPR 2022) 1.97 12.17 0.154 90.47 1.81 1.82 12.09 0.087 94.59 1.76
HPMDubbing (Cong et al., 2023) (CVPR 2023) 7.85 7.19 0.536 16.05 2.16 3.98 9.50 0.187 29.82 2.01
StyleDubber (Cong et al., 2024b) (ACL 2024) 3.87 10.92 0.607 13.14 3.02 3.74 11.00 0.540 14.18 3.04
Speaker2Dubber (Zhang et al., 2024) (MM 2024) 3.76 10.56 0.663 16.98 3.61 3.45 11.17 0.583 18.10 3.64
M2CI-Dub (Zhao et al., 2025) (ICASSP 2025) 7.99 6.91 0.621 12.85 3.15 - - - - -
EmoDubber (Cong et al., 2024a) (CVPR 2025) 8.11 6.92 0.718 11.72 3.82 8.09 6.96 0.625 12.81 3.75
@@
Table 2. The zero shot results under Dub 3.0 setting, which use unseen speaker as refernce audio.
@@
Speaker2Dubber (Zhang et al., 2024) 4.83 10.39 15.91 3.53
@@ 4.1. Implementation Details
The semantic tokenizer consists of 12 ConvNeXt blocks and 2 downsampling blocks. The codebook size of VQ is 8192. The ECAPA-TDNN in the global tokenizer features an embedding dimension of 512, while the GE2E in style flow matching prediction is used to extract speaker embedding with dimension of 256. We follow the Qwen2.5 tokenizer to process raw text. The cross model transformer is consists of 8 layers with 2 heads, and the dimension size is 256. The solver of conditional flow matching is euler, and we set the number of ODE steps is 10. In dual contrastive aligning, we use 4 heads for multi-head attention with 256 hidden sizes to obtain the attention similarity matrix. The temperature coefficient $\tau$ of $\mathcal{L}_{pm}$ and $\mathcal{L}_{mp}$ as both 0.1. We remove the vocoder bias denoiser in HiFiGAN. In data process, the video frames are sampled at 25 FPS and all audios are resampled to 16kHz. The lip region is resized to 96 $\times$ 96 and pre-trained on ResNet-18, following (Martinez et al., 2020; Ma et al., 2021). The window length, frame size, and hop length in STFT are 640, 1,024, and 160, respectively. For LLM-based Voice Enhancement Guidance, the guidance scale is set between 0.0 and 0.8 empirically. We set the batch size to 16 on Chem dataset and 64 on GRID. Our model is implemented in PyTorch. Both training and inference are implemented with PyTorch on a GeForce RTX 4090 GPU.
@@ 4.2. Datasets
We choose a real-person dubbing dataset to conduct extensive experiments to reasonably evaluate lip-sync. Our dataset mainly includes Chem and GRID.
Chem is a popular dubbing dataset recording a chemistry teacher speaking in the class (Prajwal et al., 2020a). It is collected from YouTube, with a total video length of approximately nine hours. For complete dubbing, each video has clip to sentence-level (Hu et al., 2021). The number of train, validation, and test data are 6,082, 50, and 196, respectively.
GRID is a dubbing benchmark for multi-speaker dubbing (Cooke et al., 2006). The whole dataset has 33 speakers, each with 1,000 short English samples. All participants are recorded in studio with unified background. The number of train and test data are 32,670 and 3,280.
@@ 4.3. Evaluation Metrics
We adopt multiple metrics to comprehensively evaluate the acoustic quality of synthesized dubbing. We abandon some old evaluation metrics and follow the latest speech synthesis technology to evaluate the synthesis quality. We use LSE-C/D instead of MCD-DTW-SL to evaluate lip-sync. We use SIM-O instead of SECS to evaluate speaker similarity. We adopt UTMOS instead of MCD-DTW to evaluate quality of speech. Below is the details of each metrics:
LSE-C and LSE-D. To evaluate the synchronization between the generated speech and the video quantitatively, we adopt Lip Sync Error Distance (LSE-D) and Lip Sync Error Confidence (LSE-C) as our metrics, which are widely used to lip reading (Yemini et al., 2024), talking face (Wang et al., 2023b; Jang et al., 2024), and video dubbing task (Hu et al., 2021; Lu et al., 2022). These metrics are based on the pre-trained SyncNet (Chung and Zisserman, 2016), which can explicitly test for lip synchronization in unconstrained videos in the wild (Prajwal et al., 2020b; Chung and Zisserman, 2016). Compared to the length metric MCD-SL (Chen et al., 2022), we believe that LSE-C and LSE-D can more accurately measure the synchronization of vision and audio. The discussion of the two kinds of metrics is in Appendix.
SIM-O. To evaluate the timbre consistency between the generated dubbing and the reference audio, we employ the SIM-O following (Ju et al., 2024) to compute the similarity of speaker identity. The similarity score predicted by WavLMTDNN is in the range of [-1; 1], where a larger value indicates a higher similarity of input samples.
UTMOS. UTMOS (Saeki et al., 2022) is an authoritative acoustic predictor, which focuses on evaluating the acoustic quality of synthesized speech (Zhang et al., 2025a; Ju et al., 2024; Wang et al., 2024; Du et al., 2024b; Ye et al., 2025; Wang et al., 2025), particularly by assessing naturalness, intelligibility, prosody, and expressiveness.
DNSMOS. Deep Noise Suppression MOS (DNSMOS) (Reddy et al., 2021) is designed to assess the quality of speech processed by noise suppression algorithms, measuring clarity, naturalness, background noise quality, and overall quality.
@@
Imagin (Lee et al., 2023a) (ICASSP 2023) 4.69 10.14 0.424 44.37 2.53 4.55 10.27 0.458 39.15 2.48
V2C-Net (Chen et al., 2022) (CVPR 2022) 5.59 9.52 0.430 47.82 2.41 5.34 9.76 0.356 49.09 2.40
HPMDubbing (Cong et al., 2023) (CVPR 2023) 5.76 9.13 0.461 45.51 2.14 5.82 9.10 0.359 44.15 2.11
@@
Speaker2Dubber (Zhang et al., 2024) (MM 2024) 5.27 9.84 0.734 17.04 3.69 5.19 9.93 0.606 17.00 3.73
EmoDubber (Cong et al., 2024a) (CVPR 2025) 7.12 6.82 0.802 18.53 3.83 7.10 6.89 0.665 19.75 3.81
@@
Table 4. The Clarity performance of using different scale $\alpha$ in acoustics flow matching guidance. Note that DNSMOS, SNR Score, and UTMOS are not subjective metrics.

### v2

@@ 3.3. Dual Contrastive Aligning for Dubbing
This module is designed to achieve mutual alignment between lip movement sequence and phoneme sequence by introducing a dual contrastive learning after LLM-based Semantic-aware Learning.
Lip-motion Feature Extractor. To ensure fairness for measuring alignment, we first use the same extractor (Cong et al., 2023) to obtain lip motion features from silent videos $V_{s}$:
@@
where $z_{m}\in\mathbb{R}^{{L_{v}}\times{d_{m}}}$ denotes the output lip motion embedding, $L_{v}$ indicates the length of lip sequence, and $d_{m}$ is embedding size. The $\mathrm{LipCrop}(\cdot)$ uses the face landmarks tool to crop mouth area, and $\mathrm{LipEncoder}(\cdot)$ represents the lip encoder.
Dual Contrastive Learning. We focus on learning the intrinsic correlation between phoneme-level pronunciation and lip movement to achieve reasonable alignment for movie dubbing. Following the contrastive learning manner, we introduce the InfoNCE loss (Tian, 2022) to encourage the model to distinguish correct lip-phoneme pairs. Specifically, we first treat the lip motion features $z_{m}$ as queries and the phoneme embeddings $z_{p}$ as keys. To establish positive pairs, we align each lip motion frame with its corresponding phoneme based on ground-truth timing annotations by Montreal Forced Aligner (McAuliffe et al., 2017) (MFA) and Frames Per Second (FPS). This ensures that each $z_{m}^{i}$ should be maximally similar to its temporally aligned $z_{p}^{j}$, while being distinct from other phonemes:
@@
unlike DLCL in Emodub (Cong et al., 2024a), which focuses on aligning prosody sequences (obtained by prosody adaptor) to the other (lip), our method emphasizes aligning manner between original phoneme sequences and lip to reduce the impact of prosody changes. Besides, different from the single-direction aligning in (Li et al., 2025a), our method focuses on a mutual aligning manner and does not rely on an extra duration predictor that learn coarse-grained time relevance by additional MSE loss. Finally, we use the average of mutual aligning results as dual contrastive loss:
@@
Aligning Phoneme Level Feature. The similarity matrix between phoneme embedding and lip motion embedding ${Sim}(z_{m},z_{p})$ is constrained by dual contrastive learning, then ${Sim}(z_{m},z_{p})$ further guides the hybrid generation of aligned sequences, including: (1) lip-related aligning sequences $C_{lip}$. (2) phoneme related aligning sequences. Specifically, $C_{lip}$ is obtained by multi-head attention module in  (Cong et al., 2023), in which the $z_{p}$ serves as key and value, and the $z_{m}$ is the query. Unlike (Cong et al., 2023), the learnable ${Sim}(z_{m},z_{p})$ is used as multi-head attention weight matrix to provide correct relevance. Next, by monotonic alignment search (MAS) (Kim et al., 2020), the ${Sim}(z_{m},z_{p})\in\mathbb{R}^{{L_{v}}\times{L_{t}}}$ is flat to mapping table $tab\in\mathbb{R}^{{L_{t}}\times{1}}$, which records the number of video frames corresponding to each phoneme unit. Finally, the tabtab, $LLM_{p}$, $z_{p}$, and $C_{lip}$ are associated to mel-spectrograms level prior conditions $\mu$:
@@
where $\mathrm{Up}(\cdot)$ is used to expand $LLM_{p}$ and $z_{p}$ to video level according to mapping tabtab. The $\mathcal{F}(\cdot)$ indicates the fusion module, which consists of 2D upsampling convolutional layers and transformer-based mel-decoder. The output $\mu\in\mathbb{R}^{{L_{m}}\times{d_{m}}}$, where $l_{m}$} and $d_{m}$ represent the length and embedding size of the target mel-spectrogram.
@@ 3.4. Flow-based Voice Enhancing
In this section, we introduce flow-based voice enhancing, including Style Flow Matching Prediction to inject speaker style into flow matching and LLM-based Acoustics Flow Matching Guidance to improve the clarity of generated speech via decoupled Classifier-Free Guidance (CFG) enhancement.
Table 1. Compared with related Dubbing methods on Chem benchmark. For the Dub 1.0 setting, we use the ground truth audio as reference audio, for the Dub 2.0 setting, we use the non-ground truth audio from the same speaker within the dataset as the reference audio, which is more aligned with practical usage in dubbing. $\uparrow(\downarrow)$ means that higher (lower) value is better.
@@
StyleDubber (Cong et al., 2024b) (ACL 2024) 3.87 10.92 0.607 13.14 3.14 3.74 11.00 0.501 14.18 3.04
Speaker2Dubber (Zhang et al., 2024b) (MM 2024) 3.76 10.56 0.663 16.98 3.61 3.45 11.17 0.583 18.10 3.64
@@
Table 2. The zero shot results under dubbing setting 3.0, which use unseen speaker as refernce audio.
@@
Speaker2Dubber (Zhang et al., 2024b) 4.83 10.39 15.91 3.53
@@
Style Flow Matching Prediction. Flow matching generates mel-spectrograms $\hat{M}$ from Gaussian noise by a vector field. Given mel-spectrogram space with data $M$, where $M\sim q(M)$. We aim to train a flow matching network to fit $q(M)$ by predicting the probability density path given the vector field, which can be defined as $p_{t}(x)$. Here $t\in[0,1]$, $p_{0}(x)=\mathcal{N}(x;0,I)$ and $p_{1}(x)=q(x)$. Flow matching can predict the probability density path, gradually transforming $x_{0}\sim p_{0}(x)$ into $M\sim q(M)$. Our flow matching prediction network is based on optimal-transport conditional flow matching (OT-CFM). OT-CFM uses a linear interpolation flow $\phi_{t}(x)=(1-(1-\sigma_{\min})t)x_{0}+tM$, which satisfies the marginal condition $\phi_{0}(x)=x_{0}$ and $\phi_{1}(x)=M$. The gradient field vector field of OT-CFM is $u_{t}(\phi_{t}(x)|M)=M-(1-\sigma_{min})x_{0}$. The training objective of flow matching prediction network is to predict the gradient vector field $v_{t}(\phi_{t}(x)|\mu_{SATL},\theta)$, which should be close to $u_{t}(\phi_{t}(x)|M)$: Here $\mu_{SATL}$ is style-enhanced mel-spectrogram level prior according to $\mu$ in Eq. 10. To enhance speakers’ style, we introduced SATL in flow matching. Specifically, during the flow matching generation process, SATL introduces and enhances style information through an affine transformation, which can be formulated as:
(11) $\mu_{SATL}=\gamma_{2}(\gamma_{1}\mu+\beta_{1})+\beta_{2},$
where $\gamma_{1},\gamma_{2},\beta_{1}$, and $\beta_{2}$ are parameters predicted by SATL based on style features. We train the Style Flow Matching Prediction Network using the condition $\mu_{SATL}$. We aim for the Flow Matching prediction network to generate the target mel-spectrogram M conditioned on a given $\mu_{SATL}$. During the inference process, the prediction network solves the ODE $d\phi_{t}(x)=v_{t}(\phi_{t}(x)|\mu_{SATL},\theta)dt$ from $t=0$ to $t=1$ to generate a mel-spectrogram $\hat{M}$.
LLM-based Acoustics Flow Matching Guidance. To enhance the clarity of the generated result, we enhanced the mel-spectrograms level prior conditions by LLM-based Acoustics Flow Matching Guidance. We observed that the generation process in LLM includes semantic tokens and text tokens, which introduce semantic knowledge. Specifically, we enhance LLM’s information in flow matching process to improve speech clarity based on classifier-free guidance (CFG), which can be formulated as:
(12) $\tilde{v}_{t}(\phi_{t}(x)|\mu,\theta)$ $=v_{t}(\phi_{t}(x)|\mu_{SATL},\theta)$
$+\alpha\Big{(}v_{t}(\phi_{t}(x)|\mu_{SATL},\theta)-v_{t}(\phi_{t}(x)|\mu^{\prime},\theta)\Big{)},$
where $\mu^{\prime}=\mathcal{F}([C_{lip},\mathrm{Up}(\phi,z_{p},tab))$, and $\phi$ refers to zero vector. For adapting dubbing scenarios, our flow matching explicitly decouples the condition inputs into two distinct streams: LLM-based semantic features and original features (aligning with lip movement) to improve dubbing clarity without disturbing the lip aligning prior. As a result, we can enhance only the LLM information with classifier-free guidance by controlling the scale factor $\alpha$. In general, the proposed guidance mechanism integrates LLM features as high-level semantic conditions to flow-matching network, thereby refining the gradient vector field generation process to ensure clarity while preserving the temporal correlation for audio-visual alignment.
@@ 4.1. Implementation Details
Following the Spark-TTS (Wang et al., 2025), the semantic tokenizer consists of 12 ConvNeXt blocks and 2 downsampling blocks. The codebook size of VQ is 8192. The ECAPA-TDNN in the global tokenizer features an embedding dimension of 512. The cross-modal transformer consists of 8 layers with 2 heads, and the dimension size is 256. In dual contrastive aligning, we use 4 heads for multi-head attention with 256 hidden sizes to obtain the attention similarity matrix. The temperature coefficient $\tau$ of $\mathcal{L}_{pm}$ and $\mathcal{L}_{mp}$ as both 0.1. In data processing, the video frames are sampled at 25 FPS, and all audios are resampled to 16kHz. The lip region is resized to 96 $\times$ 96 and pre-trained on ResNet-18, following (Martinez et al., 2020; Ma et al., 2021). The window length, frame size, and hop length in STFT are 640, 1,024, and 160, respectively. For LLM-based Acoustics Flow Matching Guidance, the guidance scale is set between 0.0 and 0.8 empirically. We set the batch size to 16 on Chem dataset and 64 on GRID. Both training and inference are implemented with PyTorch on a GeForce RTX 4090.
@@ 4.2. Datasets
Chem is a real-person dubbing dataset recording a chemistry teacher speaking in the class (Prajwal et al., 2020). It is collected from YouTube, with a total video length of approximately nine hours. For complete dubbing, each video has clip to sentence-level (Hu et al., 2021).
GRID is another real-person dubbing dataset (Cooke et al., 2006). The whole dataset has 33 speakers, each with 1,000 short English samples. All participants are recorded in studio with unified background.
@@ 4.3. Evaluation Metrics
We abandon some old evaluation metrics and follow the latest speech synthesis technology to evaluate the synthesis quality. Specifically, we use LSE-C/D instead of MCD-DTW-SL to evaluate lip-sync. We use SIM-O instead of SECS to evaluate speaker similarity. We adopt UTMOS instead of MCD-DTW to evaluate quality of speech. Below are the details of each metric:
LSE-C and LSE-D. Compared to the length metric MCD-DTW-SL (Chen et al., 2022), we believe that Lip Sync Error Distance (LSE-D) and Lip Sync Error Confidence (LSE-C) (Chung and Zisserman, 2016) can more accurately measure the synchronization of vision and audio. These metrics are based on the pre-trained SyncNet (Chung and Zisserman, 2016), which is widely used for lip reading (Yemini et al., 2024), talking face (Wang et al., 2023b; Jang et al., 2024), and the video dubbing task (Hu et al., 2021; Lu et al., 2022).
SIM-O. To evaluate the timbre consistency between the generated dubbing and the reference audio, we employ the SIM-O following (Ju et al., 2024) to compute the similarity of speaker identity.
UTMOS. UTMOS (Saeki et al., 2022) focuses on evaluating the acoustic quality of synthesized speech (Zhang et al., 2025a; Ju et al., 2024; Wang et al., 2024; Du et al., 2024b; Ye et al., 2025b; Wang et al., 2025), particularly by assessing naturalness, intelligibility, prosody, and expressiveness.
DNSMOS. Deep Noise Suppression MOS (DNSMOS) (Reddy et al., 2021) is designed to assess the quality of speech processed by noise suppression algorithms, measuring clarity.
@@
Speaker2Dubber (Zhang et al., 2024b) (MM 2024) 5.27 9.84 0.734 17.04 3.69 5.19 9.93 0.606 17.00 3.73
@@
Table 4. The Clarity performance of using different scale $\alpha$ in acoustics flow matching guidance. Note that DNSMOS, SNR Score, and UTMOS are not human subjective metrics.